OpenStudy (anonymous):

Given that P=(-7,16) and Q=(-8,7), find the component form and magnitude of QP PLEASE HELP MEDAL AND FOLLOW WILL BE GIVEN!!! @sleepyhead314

jimthompson5910 (jim_thompson5910):

"find the component form" you subtract the corresponding coordinates in the same order (tail - head) y2 - y1 = 7 - 16 = -9 x2 - x1 = -8 - (-7) = -8 + 7 = -1 so PQ in component form is <-1, -9> this tells us that to go from (-7,16) to (-8,7) we need to go 1 unit to the left and 9 units down

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

all I do is subtract the ys and the xs? why didn't some one tell me that earlier! is the answer <-1,-9>, square root 82?

jimthompson5910 (jim_thompson5910):

yep you just subtract the ys to get the change in y subtract the xs to get the change in x the order is (tail - head) which means you start with the ending point and subtract off the starting point

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{82}\] is the correct answer to the second part

jimthompson5910 (jim_thompson5910):

this is the distance from P to Q, in other words, the length of the vector

OpenStudy (anonymous):

thank you can you help me with a couple more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Evaluate the expression r=<7,-3,-7> V=<4,6,-5> w=<-5,-6,-3> v-w

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

v - w is the same idea as finding the component form you subtract the coordinates (start with v, subtract off w)

jimthompson5910 (jim_thompson5910):

example: the x coordinate of v is x = 4, the x coordinate of w = -5 subtract the x coords: Vx - Wx = 4 - (-5) = 4 + 5 = 9

OpenStudy (anonymous):

I did but I didn't get anything close to my options

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

and my options are -41 <-20,-36,15> <-28,18,-35> -9

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you sure it's v - w? and it's not like 2v - 3w ?

OpenStudy (anonymous):

yes it just says v-w

jimthompson5910 (jim_thompson5910):

well that's just odd

jimthompson5910 (jim_thompson5910):

one of the options should have 9 in the x coordinate

OpenStudy (anonymous):

which one is the closest?

jimthompson5910 (jim_thompson5910):

none of them sadly, idk what to do here. Can you post a screenshot of the full thing?

OpenStudy (anonymous):

no but ill retype it evaluate the expression r=<7,-3,-7> v=<4,6,-5> w=<-5,-6,-3> v-w

OpenStudy (anonymous):

that's all it gives me

jimthompson5910 (jim_thompson5910):

it's also odd how they mention vector r but don't use it in v-w

jimthompson5910 (jim_thompson5910):

it seems like there is a typo, but idk

OpenStudy (anonymous):

me ether -.- thanks though

jimthompson5910 (jim_thompson5910):

sadly there's not much I can do other than tell you to ask your teacher about it. S/he would know if there's a typo or not. If so, then you'd get free points